# Reblurring-Guided Single Image Defocus Deblurring: A Learning Framework with Misaligned Training Pairs

### v1

@@ Abstract
For single image defocus deblurring, acquiring well-aligned training pairs (or training triplets), i.e., a defocus blurry image, an all-in-focus sharp image (and a defocus blur map), is an intricate task for the development of deblurring models. Existing image defocus deblurring methods typically rely on training data collected by specialized imaging equipment, presupposing that these pairs or triplets are perfectly aligned. However, in practical scenarios involving the collection of real-world data, direct acquisition of training triplets is infeasible, and training pairs inevitably encounter spatial misalignment issues. In this work, we introduce a reblurring-guided learning framework for single image defocus deblurring, enabling the learning of a deblurring network even with misaligned training pairs. Specifically, we first propose a baseline defocus deblurring network that utilizes spatially varying defocus blur map as degradation prior to enhance the deblurring performance. Then, to effectively learn the baseline defocus deblurring network with misaligned training pairs, our reblurring module ensures spatial consistency between the deblurred image, the reblurred image and the input blurry image by reconstructing spatially variant isotropic blur kernels. Moreover, the spatially variant blur derived from the reblurring module can serve as pseudo supervision for defocus blur map during training, interestingly transforming training pairs into training triplets. Additionally, we have collected a new dataset specifically for single image defocus deblurring (SDD) with typical misalignments, which not only substantiates our proposed method but also serves as a benchmark for future research. The effectiveness of our method is validated by notable improvements in both quantitative metrics and visual quality across several datasets with real-world defocus blurry images, including DPDD, RealDOF, DED, and our SDD. The source code and dataset are available at https://github.com/ssscrystal/Reblurring-guided-JDRL.
@@ 1 Introduction
However, for other consumer cameras, e.g., digital single lens reflex or smartphone cameras, the availability of training triplets is impractical. In widely used datasets for image defocus deblurring, such as DPDD [12] and RealDOF [14] datasets, meticulous control is exercised over the capturing camera to ensure consistent acquisition of training pairs comprising a defocus blurry image and a ground-truth sharp image. Nevertheless, as illustrated in Fig. 1, inevitable misalignments still occur. We also note that different imaging sensors introduce various types of defocus blur, posing challenges for learned deblurring models based on specific cameras when handling cases involving other sensor types, referring to Fig. 11. Therefore, effective deployment on target devices requires learning a device-specific image defocus deblurring model while relaxing the requirement for perfect alignment in training pairs.
@@
In this work, we propose a novel learning framework for single image defocus deblurring, as shown in Fig. 2, specifically focusing on effectively learning a single image defocus deblurring model from misaligned training pairs that can be easily obtained for a given imaging camera. We first introduce a baseline image defocus deblurring model that utilizes a lightweight blur prediction subnetwork to provide degradation prior and guide the deblurring process. Then, we propose a reblurring-based learning framework, as shown in Fig. 5, to address the challenge of misaligned training pairs. To ensure the consistent spatial alignment between the deblurred image, the reblurred image and the input blurry image, our reblurring module consists of a kernel prediction network and weight prediction network to reconstruct spatially variant isotropic blur kernels. Moreover, pseudo defocus blur map can be derived from the reblurring module, and is then utilized to guide the training of defocus blur prediction subnetwork, thereby expanding the training pairs to training triplets.
@@
Furthermore, we introduce a new dataset named SDD for single image defocus deblurring. The image pairs are captured using a HUAWEI X2381-VG camera, which can be adjusted to capture pairs of blurry and sharp images by manipulating the camera motor or aperture size. Despite our efforts to maintain alignment during the collection process of the SDD dataset, some misalignment persists due to variations in consumer-grade cameras and collection settings. The misalignment primarily manifests in two forms: zoom misalignment and shift misalignment, as illustrated in Fig. 1. Importantly, the degree of misalignment within the SDD dataset tends to be more severe compared to that observed in the DPDD dataset, making it be a testbed for evaluating our reblurring-guided image defocus deblurring techniques and serve as a benchmark for future research in this field.
@@
This work is previously presented as a conference paper with oral presentation [21], upon which this manuscript has made three major improvements: (i) We have enhanced the learning framework by enabling the derivation of pseudo defocus blur map from the reblurring module and constructing triplets in training datasets. (ii) Correspondingly, we have improved the deblurring network instead of directly borrowing existing works such as UNet and MPRNet, as done in [21]. (iii) To provide a comprehensive comparison, we have included additional state-of-the-art methods for evaluation, including Restormer [15], DID-ANet [19], Loformer [22] and GRLS-B [18]. Furthermore, we have incorporated the DED dataset [19] to evaluate competing methods.
In summary, the contributions of this paper include
@@
    A novel reblurring-guided learning framework for image defocus deblurring that effectively exploits misaligned training pairs for learning deblurring models,
@@
    A baseline defocus deblurring network based on training triples, where pseudo defocus map supervision can be derived from our reblurring module, and
@@
    A new dataset named SDD comprising high-resolution image pairs with diverse contents for evaluating image defocus deblurring models and facilitating future research in this field.
@@ 2.1 Defocus Deblurring Methods
Recent approaches predominantly utilize deep learning to overcome the limitations of the traditional methods by restoring the image directly from the blurred image. Abuolaim et al. [12] pioneered the first end-to-end learning-based method DPDNet. This approach achieved significantly better results than traditional two-stage methods, establishing a new benchmark in the field. Subsequently, Lee et al. [14] designed a network featuring an iterative filter adaptive module to address spatially varying defocus blur. Son et al. [24] proposed a kernel prediction adaptive convolution technique that further refines the capacity to address complex defocus patterns. Despite these advancements, it is crucial to note that the effectiveness of these deep learning-based defcus deblurring methods heavily depends on the quality of the training data, which is primarily derived from the DPDD dataset. The dependence on high-quality, well-aligned training pairs has been a persistent challenge, motivating our research to focus on single image defocus deblurring with misaligned training data. Although some weakly supervised or unsupervised methods[25, 26, 27] can partially address the reliance on large-scale data, these methods often do not perform as well as supervised approaches.
@@ 2.3 Reblurring Process
Figure 2: Overview of reblurring-guided learning framework for image defocus deblurring. It consists of deblurring network $\mathcal{F}$ and reblurring network $\mathcal{R}$. In $\mathcal{F}$, defocus blur map is integrated as degradation prior for enhancing deblurring performance. To tackle spatial misalignment when training $\mathcal{F}$, optical flow-based deformation $\mathcal{W}$ is adopted to accommodate misalignment, while reblurring network $\mathcal{R}$ ensures spatial consistency of the deblurred image, the reblurred image and the input blurry image. Pseudo supervision of defocus blur map can be derived from reblurring network.
@@ 3 Proposed Method
Recent advancements in image defocus deblurring have primarily focused on the development of learning-based models using training pairs, denoted as $\{\bm{I}_{B}^{n},\bm{I}_{S}^{n}\}_{n=1}^{N}$, where $\bm{I}_{B}$ represents a defocus blurry image and $\bm{I}_{S}$ denotes a ground-truth sharp image. However, even with careful alignment during data collection and post-processing techniques, spatial misalignment remains an inevitable issue as depicted in Fig. 1. Furthermore, when considering practical applications with new sensors, the severity of spatial misalignment issues may increase compared to those encountered in DPDD. Additionally, it is important to note that different imaging sensors exhibit distinct patterns of defocus blur which limits the generalization ability of learned deblurring models to real-world cases captured by other devices. Therefore, for effective deployment on target devices, it is necessary to learn a image defocus deblurring model specifically tailored for each device while relaxing the requirement for perfect alignment in training pairs.
@@
To sum up, learning deblurring models based on misaligned training pairs is both challenging and meaningful.
In the following, we first provide an overview of our proposed reblurring-based learning framework, which effectively leverages misaligned training pairs to learn image defocus deblurring models. Subsequently, we offer detailed explanations on the baseline deblurring model, reblurring process, derivation of pseudo defocus blur map, and finally introduce our newly established dataset.
@@ 3.1 The Overall Framework
      Input       UNet       Ours       GT
Figure 3: The direct utilization of pixel-wise loss function for training a UNet model based on misaligned training pairs often leads to deformation artifacts, such as distorted brick lines. These artifacts can be effectively alleviated by employing our reblurring-based learning framework for training a deblurring model. In order to ensure a fair comparison in this case, we utilize the same Unet architecture as our deblurring model, and training loss is Eq. (1).
Given a training set with $N$ pairs $\{\bm{I}_{B}^{n},\bm{I}_{S}^{n}\}_{n=1}^{N}$, where the blurry image $\bm{I}_{B}$ is not perfectly aligned with the ground-truth sharp image $\bm{I}_{S}$, our objective is to learn a deblurring model that effectively addresses misalignment issues while minimizing deformation distortions. In this work, we propose a novel reblurring-guided learning framework, wherein the misalignment problem can be resolved through the integration of a reblurring module. As illustrated in Fig. 2, our proposed framework comprises two main components: a deblurring module, and a reblurring module. The deblurring module utilizes a dedicated network $\mathcal{F}$ to process the input blurry image $\bm{I}_{B}$ and generate an estimated deblurred image $\hat{\bm{I}}$. To address misalignment between the ground-truth sharp image $\bm{I}_{S}$ and its corresponding estimate $\hat{\bm{I}}$, we introduce an optical flow-based deformation strategy integrated in our approach instead of relying solely on direct pixel-wise loss. In this manner, the deblurred result $\hat{\bm{I}}$ can adaptively learn sharp textures from the sharp image $\bm{I}_{S}$, thereby liberating it from the constraints of pixel-level precision. To address potential artifacts caused by optical flow deformation, we introduce a calibration mask and cycle deformation, which are further elaborated in Section 3.2. Subsequently, the reblurring module ensures spatial consistency between $\hat{\bm{I}}$ and $\bm{I}_{B}$. For this purpose, we propose a reblurring network $\mathcal{R}$ tasked with generating a reblurred image $\hat{\bm{I}}_{B}$ that closely approximates $\bm{I}_{B}$. Meanwhile, to ensure spatial coherence between $\hat{\bm{I}}$ and $\bm{I}_{B}$, the reblurring network $\mathcal{R}$ can also predict the isotropic blur kernels in polar coordinates.
The training loss for learning the parameters in deblurring network $\mathcal{F}$ and reblurring network $\mathcal{R}$ can be formally expressed as
where $\mathcal{L}_{d}$ is deblurring loss, $\mathcal{L}_{r}$ is reblurring loss, and $\alpha$ is a trade-off parameter. Benefiting from the reblurring-guided training strategy, a Unet can be trained to be free from deformation artifacts, as shown in Fig. 3.
In this work, we further suggest that incorporating a defocus blur map can enhance the deblurring performance of our deblurring model. Specifically, our model first estimates the defocus map and then utilizes it to improve the deblurring effect. The utilization of degradation-related priors as input has been validated in various tasks  [19, 42], as estimating the degradation is comparatively easier than performing deblurring itself. To enable the training of our baseline deblurring model, we need to prepare training triples, denoted as $\{\bm{I}_{B}^{n},\bm{I}_{S}^{n},\bm{B}^{n}\}_{n=1}^{N}$, where $\bm{B}$ represents the ground-truth spatially variant defocus blur map. However, obtaining such ground-truth maps is impractical due to difficulties in capturing spatially variant blur accurately. While previous work  [19] estimated defocus blur maps from light field data captured by a Lytro camera, this approach is not suitable for popular consumer cameras. Fortunately, our proposed reblurring module provides a means to obtain pseudo ground-truth defocus maps $\bm{B}_{P}$ easily for each input blurry image and serves as supervision for the sub-network within our deblurring model. Therefore, based on the training triplets $\{\bm{I}_{B}^{n},\bm{I}_{S}^{n},\bm{B}_{P}^{n}\}_{n=1}^{N}$, the final training loss can be defined as
where $\mathcal{L}_{P}$ is a loss function for defocus blur map, and $\beta$ is a positive trade-off parameter.
In this way, we have developed a robust framework that enables the deblurring network $\mathcal{F}$ to effectively leverage spatially adaptive sharp textures from the ground-truth sharp image while maintaining spatial consistency with the blurry input. During inference, our baseline deblurring model can generate latent sharp images while discarding the reblurring module.
3.2 Deblurring Module
In the deblurring module, we introduce a baseline deblurring model $\mathcal{F}$ along with its training loss $\mathcal{L}_{d}$. By leveraging this loss function $\mathcal{L}_{d}$, the deblurred output $\hat{\bm{I}}$ obtained from $\mathcal{F}$ can dynamically acquire high-resolution textures from the sharp image $\bm{I}_{S}$, thereby transcending the limitations imposed by pixel-level precision.
[…]
Figure 4: Errors in optical flow estimation can incorrectly deform the sharp ground-truth image, e.g. in areas marked by green rectangles. Calibration masks can help by filtering out this adverse region.
@@ 3.2 Deblurring Module
3.2.1 Deblurring Model $\mathcal{F}$
Our baseline deblurring model $\mathcal{F}$ takes the blurry image $\bm{I}_{B}$ as input and generates an intermediate defocus blur map $\hat{\bm{B}}$ and a latent sharp image $\hat{\bm{I}}$. The estimated defocus blur map, denoted as $\hat{\bm{B}}$, serves as a degradation prior to enhance the deblurring performance within the deblurring network. Specifically, our baseline model $\mathcal{F}$ consists of two subnetworks: defocus blur map prediction $\mathcal{F}_{B}$ and deblurring sub-network $\mathcal{F}_{D}$, which collectively produce the final deblurred output
$\hat{\bm{I}}=\mathcal{F}_{D}(\bm{I}_{B},\mathcal{F}_{B}({\bm{I}_{B}})).$ (3)
The defocus map prediction $\mathcal{F}_{B}$ is a lightweight network composed of several convolutional layers and residual layers, which can effectively learn from the pseudo ground-truth defocus maps provided by our reblurring module. The defocus map and the blurred image are fused through a Channel Attention mechanism which dynamically adjusts the importance of each channel by re-weighting the channels of the defocus maps and the blurred images. This allows the model to better focus on the important features related to the degree of defocus blur, thereby enhancing the performance of the deblurring network. The processed defocus map and blurred image are then fed into $\mathcal{F}_{B}$ composed of transformer blocks for deblurring. The deblurring $\mathcal{F}_{D}$ has a 4-level symmetric encoder-decoder as [15]. The parameters of $\mathcal{F}_{D}$ and $\mathcal{F}_{B}$ can be learned by optimizing a single deblurring loss $\mathcal{L}_{D}$. However, this approach does not provide supervision for the estimated defocus blur map $\hat{\bm{B}}$, which may result in significant deviation from the true defocus blur map. To address this issue, we introduce an additional loss term $\mathcal{L}_{P}$ based on the pseudo defocus blur map $\bm{B}_{P}$, as explained in Section 3.4. In this section, we mainly focus on the deblurring loss function $\mathcal{L}_{D}$.
3.2.2 Deblurring Loss $\mathcal{L}_{D}$
where $\mathcal{W}$ denotes linear warping operation [43].
3.3 Reblurring Module
Following the deblurring process, the reblurring module is employed to enhance the spatial consistency of the deblurred image. The objective of the reblurring module $\mathcal{R}$ is to maintain the spatial coherence between $\hat{\bm{I}}$ and $\bm{I}_{B}$. A spatially invariant reblurring network $\mathcal{R}$ is utilized to reintroduce blurriness into $\hat{\bm{I}}$, guided by predicted isotropic blurring kernels. Subsequently, pixel-wise loss is applied to supervise $\hat{\bm{I}}_{B}$ with respect to its correspondence with $\bm{I}_{B}$.
@@ 3.3 Reblurring Module
3.3.1 Reblurring Network $\mathcal{R}$

### v2

@@ Abstract
For single image defocus deblurring, acquiring well-aligned training pairs (or training triplets), i.e., a defocus blurry image, an all-in-focus sharp image (and a defocus blur map), is a challenging task for developing effective deblurring models. Existing image defocus deblurring methods typically rely on training data collected by specialized imaging equipment, with the assumption that these pairs or triplets are perfectly aligned. However, in practical scenarios involving the collection of real-world data, direct acquisition of training triplets is infeasible, and training pairs inevitably encounter spatial misalignment issues. In this work, we introduce a reblurring-guided learning framework for single image defocus deblurring, enabling the learning of a deblurring network even with misaligned training pairs. By reconstructing spatially variant isotropic blur kernels, our reblurring module ensures spatial consistency between the deblurred image, the reblurred image and the input blurry image, thereby addressing the misalignment issue while effectively extracting sharp textures from the all-in-focus sharp image. Moreover, spatially variant blur can be derived from the reblurring module, and serve as pseudo supervision for defocus blur map during training, interestingly transforming training pairs into training triplets. To leverage this pseudo supervision, we propose a lightweight defocus blur estimator coupled with a fusion block, which enhances deblurring performance through seamless integration with state-of-the-art deblurring networks. Additionally, we have collected a new dataset for single image defocus deblurring (SDD) with typical misalignments, which not only validates our proposed method but also serves as a benchmark for future research. The effectiveness of our method is validated by notable improvements in both quantitative metrics and visual quality across several datasets with real-world defocus blurry images, including DPDD, RealDOF, DED, and our SDD. The source code and dataset are available at https://github.com/ssscrystal/Reblurring-guided-JDRL.
@@ 1 Introduction
However, for other consumer cameras, e.g., digital single lens reflex or smartphone cameras, the availability of training triplets is impractical. In widely used datasets for image defocus deblurring, such as DPDD [1] and RealDOF [2] datasets, meticulous control is exercised over the capturing camera to ensure consistent acquisition of training pairs comprising a defocus blurry image and a ground-truth sharp image. Nevertheless, as illustrated in Fig. 1, inevitable misalignments still occur. We also note that different imaging sensors introduce various types of defocus blur, posing challenges for learned deblurring models based on specific cameras when handling cases involving other sensor types. Therefore, effective deployment on target devices requires learning a device-specific image defocus deblurring model while relaxing the requirement for perfect alignment in training pairs.
In this work, we propose a novel learning framework for single image defocus deblurring (Fig. 2), specifically focusing on effectively learning a single image defocus deblurring model from misaligned training pairs that can be easily obtained for a given imaging camera.
@@
We first develop a reblurring-based learning framework to address the challenge of misaligned training pairs. To ensure the consistent spatial alignment between the deblurred image, the reblurred image and the input blurry image, our reblurring module consists of a Kernel Prediction Network (KPN) and a Weight Prediction Network (WPN) to reconstruct spatially variant isotropic blur kernels. Moreover, spatially variant blur can be derived from the reblurring module, and serve as pseudo supervision for defocus blur map during training, interestingly expanding the training pairs to training triplets. To employ the pseudo supervision, we design a lightweight defocus blur estimator coupled with a fusion block, which enhances deblurring performance through seamless integration with state-of-the-art deblurring networks.
@@
Furthermore, we introduce a new dataset named SDD for single image defocus deblurring. The image pairs are captured using a HUAWEI X2381-VG camera, which can be adjusted to capture pairs of blurry and sharp images by manipulating the camera motor or aperture size. Despite our efforts to maintain alignment during the collection process of the SDD dataset, some misalignment persists due to variations in consumer-grade cameras and collection settings. The misalignment primarily manifests in two forms: zoom misalignment and shift misalignment, as illustrated in Fig. 1. Importantly, the degree of misalignment within the SDD dataset tends to be more severe compared to that observed in the DPDD [1] dataset, making it be a testbed for evaluating our reblurring-guided image defocus deblurring techniques and serve as a benchmark for future research in this field.
@@
This work is previously presented as a conference paper with oral presentation [21], upon which this manuscript has made three major improvements: (i) We have enhanced the learning framework by deriving pseudo defocus blur maps from the reblurring module and constructing triplets in training datasets. (ii) Correspondingly, we have improved the deblurring network architecture by incorporating a lightweight defocus blur map estimator coupled with a fusion block. This design not only seamlessly integrates with existing deblurring models but also significantly enhances deblurring performance. (iii) To provide a comprehensive comparison, we have included state-of-the-art methods based on convolutional neural networks (CNN) and Transformer for evaluation, including UformerT [16], Restormer [15], DID-ANet [19] and Loformer [22]. Furthermore, we have incorporated the DED dataset [19] to evaluate competing methods.
In summary, the contributions of this paper can be summarized as
@@
    A novel reblurring-guided learning framework is proposed for image defocus deblurring that effectively exploits misaligned training pairs for learning deblurring models.
@@
    A lightweight blur map estimator and a fusion block are designed to integrate with state-of-the-art deblurring networks, where the pseudo supervision on spatially variant blur map can be derived from our reblurring module.
@@
    A new dataset named SDD comprising high-resolution image pairs with diverse contents is introduced for evaluating image defocus deblurring models and facilitating future research in this field. On benchmark datasets including DPDD [1], RealDOF [2], DED [19] and our SDD, our method is significantly superior to existing methods.
@@ 2.1 Defocus Deblurring Methods
Recent approaches predominantly utilize deep learning to overcome the limitations of the traditional methods by restoring the image directly from the blurred image. Abuolaim et al. [1] pioneered the first end-to-end learning-based method DPDNet. This approach achieved significantly better results than traditional two-stage methods, establishing a new benchmark in the field. Subsequently, Lee et al. [2] designed a network featuring an iterative filter adaptive module to address spatially varying defocus blur. Son et al. [24] proposed a kernel prediction adaptive convolution technique that further refines the capacity to address complex defocus patterns. Despite these advancements, it is crucial to note that the effectiveness of these deep learning-based defcus deblurring methods heavily depends on the quality of the training data, which is primarily derived from the DPDD [1] dataset. The dependence on high-quality, well-aligned training pairs has been a persistent challenge, motivating our research to focus on single image defocus deblurring with misaligned training data. Although some weakly supervised or unsupervised methods[25, 26, 27] can partially address the reliance on large-scale data, these methods often do not perform as well as supervised approaches.
@@ 2.2 Defocus Deblurring Datasets
Figure 2: Overview of our reblurring-guided learning framework for image defocus deblurring. It consists of deblurring module and reblurring module. In deblurring module, our introduced blur estimator $\mathcal{E}$ and fusion block $\mathcal{F}$ provide spatially variant degradation priors for enhancing deblurring performance, and can be seamlessly integrated with existing deblurring network $\mathcal{D}$. To tackle spatial misalignment of training pairs, optical flow-based deformation $\mathcal{W}$ is adopted to accommodate misalignment, while reblurring network $\mathcal{R}$ ensures spatial consistency of the deblurred image, the reblurred image and the input blurry image. Pseudo supervision $\bm{B}_{P}$ of defocus blur map $\hat{\bm{B}}$ can be derived from reblurring network.
@@ 3 Proposed Method
Recent advancements in image defocus deblurring have primarily focused on the development of learning-based models using training pairs, denoted as $\{\bm{I}_{B}^{n},\bm{I}_{S}^{n}\}_{n=1}^{N}$, where $\bm{I}_{B}$ represents a defocus blurry image and $\bm{I}_{S}$ denotes a ground-truth sharp image. However, even with careful alignment during data collection and post-processing techniques, spatial misalignment remains an inevitable issue as depicted in Fig. 1. Furthermore, when considering practical applications with new sensors, the severity of spatial misalignment issues may increase compared to those encountered in DPDD[1]. Additionally, it is important to note that different imaging sensors exhibit distinct patterns of defocus blur which limits the generalization ability of learned deblurring models to real-world cases captured by other devices. Therefore, for effective deployment on target devices, it is necessary to learn a image defocus deblurring model specifically tailored for each device while relaxing the requirement for perfect alignment in training pairs.
@@
To sum up, learning deblurring models based on misaligned training pairs is both challenging and meaningful. In the following, we first provide an overview of our proposed reblurring-based learning framework, which effectively leverages misaligned training pairs to learn image defocus deblurring models. Subsequently, we offer detailed explanations on the reblurring module with derivation of pseudo defocus blur map, the deblurring model equipped with a defocus blur map estimator and a fusion block, and finally introduce our newly established dataset.
@@ 3.1 The Overall Framework
      Input       UNet       UNet       GT
Figure 3: The direct utilization of pixel-wise loss function for training a UNet model based on misaligned training pairs often leads to deformation artifacts, such as distorted brick lines. These artifacts can be effectively alleviated by our UNet, where our reblurring-based learning framework Eq. (2) is adopted for training.
@@
Given a training set with $N$ pairs $\{\bm{I}_{B}^{n},\bm{I}_{S}^{n}\}_{n=1}^{N}$, where the blurry image $\bm{I}_{B}$ is not perfectly aligned with the ground-truth sharp image $\bm{I}_{S}$, our objective is to learn a deblurring model that effectively addresses misalignment issues while minimizing deformation distortions. In this work, we propose a novel reblurring-guided learning framework, wherein the misalignment problem can be resolved through the integration of a reblurring module.
As illustrated in Fig. 2, our proposed framework comprises two main components: a deblurring network, and a reblurring network. The deblurring netwok utilizes a dedicated network $\mathcal{D}$ to process the input blurry image $\bm{I}_{B}$ and generate an estimated deblurred image $\hat{\bm{I}}$.
To address misalignment between the ground-truth sharp image $\bm{I}_{S}$ and its corresponding estimate $\hat{\bm{I}}$, we introduce an optical flow-based deformation strategy integrated in our approach instead of relying solely on direct pixel-wise loss. In this manner, the deblurred result $\hat{\bm{I}}$ can adaptively learn sharp textures from the sharp image $\bm{I}_{S}$, thereby liberating it from the constraints of pixel-level precision. To address potential artifacts caused by optical flow deformation, we introduce a calibration mask and cycle deformation, which are further elaborated in Section 3.3. Subsequently, the reblurring module ensures spatial consistency between $\hat{\bm{I}}$ and $\bm{I}_{B}$. For this purpose, we propose a reblurring network $\mathcal{R}$ tasked with generating a reblurred image $\hat{\bm{I}}_{B}$ that closely approximates $\bm{I}_{B}$. Meanwhile, to ensure spatial coherence between $\hat{\bm{I}}$ and $\bm{I}_{B}$, the reblurring network $\mathcal{R}$ can also predict the isotropic blur kernels in polar coordinates.
The training loss for learning the parameters in deblurring network $\mathcal{D}$ and reblurring network $\mathcal{R}$ can be formally expressed as
where $\mathcal{L}_{D}$ (Eq. (14)) is deblurring loss, $\mathcal{L}_{R}$ (Eq. (7)) is reblurring loss, and $\alpha$ is a trade-off parameter. Benefiting from the reblurring-guided training strategy, a Unet can be trained to be free from deformation artifacts, as shown in Fig. 3.
In this work, we further suggest that incorporating a defocus blur map can enhance the deblurring performance of deblurring networks. Specifically, an existing deblurring network $\mathcal{D}$ can be incorporated with a defocus blur map estimator $\mathcal{E}$ coupled with a fusion block $\mathcal{F}$ to enable the utilization of predicted degradation prior for enhancement of deblurring performance. The estimator $\mathcal{E}$ predicts the defocus blur map $\hat{\bm{B}}$, while $\mathcal{F}$ incorporates the degradation prior to enhance deblurring performance using a deformable attention mechanism. The utilization of degradation-related priors as input has been validated in various tasks  [19, 42], as estimating the degradation is comparatively easier than performing deblurring itself. To enable the training of our baseline deblurring model, we need to prepare training triples, denoted as $\{\bm{I}_{B}^{n},\bm{I}_{S}^{n},\bm{B}^{n}\}_{n=1}^{N}$, where $\bm{B}$ represents the ground-truth spatially variant defocus blur map. However, obtaining such ground-truth maps is impractical due to difficulties in capturing spatially variant blur accurately. While previous work  [19] estimated defocus blur maps from light field data captured by a Lytro camera, this approach is not suitable for popular consumer cameras. Fortunately, our proposed reblurring module provides a means to obtain pseudo ground-truth defocus maps $\bm{B}_{P}$ easily for each input blurry image and serves as supervision for the defocus blur map estimator $\mathcal{E}$ within our deblurring model. Therefore, based on the training triplets $\{\bm{I}_{B}^{n},\bm{I}_{S}^{n},\bm{B}_{P}^{n}\}_{n=1}^{N}$, the final training loss can be defined as
where $\hat{\bm{B}}=\mathcal{E}({\bm{I}_{B}})$ denotes the estimated defocus blur map, $\mathcal{L}_{P}$ (Eq. (15)) is a loss function for defocus blur map, and $\beta$ is a positive trade-off parameter.
In this way, we have developed a robust framework that enables the deblurring network to effectively leverage spatially adaptive sharp textures from the ground-truth sharp image while maintaining spatial consistency with the blurry input. During inference, our baseline deblurring model can generate latent sharp images while discarding the reblurring module.
@@
3.2 Reblurring Module
In this section, We first introduce our reblurring network along with its loss function $\mathcal{L}_{R}$, and then present the acquisition of pseudo defocus blur map $\bm{B}_{P}$.
3.2.1 Reblurring Network